# The Landscape of College-level Data Visualization Courses, and the Benefits of Incorporating Statistical Thinking

### v1

@@ 4 Conclusion and Discussion
First, we conducted a survey of data visualization courses taught at 135 top colleges and universities in the United States in the 2022-2023 academic year, in order to understand the landscape of data visualization courses. We identified 270 courses across many departments; the most common were related to computer science, information systems, management, and analytics, but departments in business, the humanities, journalism, political science, the visual arts, and other fields were also represented. Thus, the instruction of data visualization is widespread and diverse. Most courses were not taught by statistics or data science departments, and thus understandably did not focus on statistical topics. Instead, most courses focused on visual storytelling, aesthetic design, dashboard design, specialized software, or other topics unique to their respective disciplines. It was especially uncommon for courses to focus on core aspects of statistical thinking: testing, uncertainty quantification, and modeling. Thus, statistics-related departments can provide a valuable perspective on data visualization that is unique to current course offerings.
@@
Second, we outlined three principles for teaching data visualization to encourage statistical thinking, using illustrative examples from our undergraduate data visualization course at Carnegie Mellon University. The first principle is that graphs should be complemented with inference. For example, the “statistical significance” of visual observations (e.g. differences in proportions or means) can be clarified with standard tests (e.g. chi-squared tests or ANOVA). This helps students understand how to combine graphs and analyses to make a cohesive argument, and encourages them to grapple with complex inferential issues (e.g. power, multiple testing, and visually overlapping confidence intervals). The second principle is that graphs are estimates and statistics that can be leveraged in an inferential framework. For example, graphs can have a null distribution that’s used in a testing procedure, and have uncertainty that can be quantified theoretically or via simulation with a bootstrap procedure. This helps students gain a deeper understanding of distributions and estimators by seeing how they operate with complex visuals beyond just univariate random variables. The third principle is that graphs can be used to motivate, teach, and interpret statistical analyses. Because visuals are a tried-and-true teaching tool, data visualization courses can be a natural conduit for students to engage with challenging concepts, like interactions in regression and principal component analysis. In this way, we’ve used data visualization as part of a spiral learning approach in our undergraduate curriculum, where students revisit topics several times (e.g. in a visualization context and in a theoretical context) as they master difficult material.
@@
That said, there are several limitations to our contributions, which motivate future work. Our survey is likely not a census of all data visualization courses, because (1) we used select keywords to identify potential courses, (2) we relied only on publicly available information to declare a course as a data visualization course, and (3) we only considered top colleges and universities according to the U.S. News and World Report in 2023. Furthermore, we focused primarily on courses’ statistical content and did not form a full taxonomy of topics taught in data visualization courses. For example, we observed that many courses focused on visual storytelling and aesthetic design (particularly in business, journalism, the visual arts, and writing), but did not study which methods within storytelling and design these courses taught. Because different disciplines have their own specialized interest in data visualization, other fields may want to study which topics are covered in data visualization courses within their own discipline, perhaps using our publicly available dataset on these courses. Indeed, we hope our dataset allows others to explore the diversity of data visualization courses, including instructors and students seeking datasets for undergraduate projects.
@@ Acknowledgments
We would like to thank Peter Freeman and Alex Reinhart for helpful comments that improved the presentation of this work.
@@
Here we discuss details about course design for our data visualization courses in the Statistics and Data Science Department at Carnegie Mellon University. Specifically, we outline the structure and content for our undergraduate and Master’s courses. Each course has its own intended audience and length. Thus, we hope these materials demonstrate how instructors in statistics-related fields can teach data visualization at the undergraduate and graduate level.
@@
Some readers may wonder how the set of students can be so diverse when the prerequisite is seemingly strict. One of our college’s general education requirements is that students must complete an introductory statistics course during their freshman year. Thus, students only have to take one additional statistics course during their undergraduate in order to enroll in our course. Our reasoning for this prerequisite is to ensure that students have some familiarity with statistical inference, such that they can appreciate how inferential tools can complement data visualizations, which is one focus of the course.
Each week consists of two 50-minute lectures (one on Monday and one on Wednesday) and a 50-minute computer lab on Friday. Typically, 1-2 lectures are dedicated to ways to visualize particular data types (e.g. two-dimensional categorical data, one-dimensional quantitative data), and then 1-2 lectures are dedicated to inferential tools that can be used for that data type (e.g. chi-squared tests, Kolmogorov-Smirnov tests). Meanwhile, in the computer labs, students work through practice problems related to that week’s lecture content; along the way, they can ask each other and the teaching staff questions, such that the computer labs act as an active learning component. The typical week-by-week timeline and set of topics for the course is shown in Table 5. The first half of the semester focuses on visualizations and inferential tools involving one or two categorical and/or quantitative variables; this also corresponds to the taxonomy illustrated in Table 1. Meanwhile, the second half of the semester focuses on topics involving more complex data types: high-dimensional, spatial, longitudinal, and text.
@@
As for assessments, we have students submit weekly homework assignments, which involve creating visualizations and conducting analyses with real datasets based on the past week’s lecture and computer lab content. Furthermore, as a small participation grade, students are asked to critique a data visualization they found online or elsewhere once a month, in order to see how principles we discuss in class apply to real-world graphics. At midsemester, students complete a take-home midterm exam that assesses their ability to implement, navigate, and interpret the visualizations and inferential tools in Table 1. After the midterm exam, students are organized into teams of 3-4 people, and by the end of the semester complete a final project that consists of visualizations and analyses of a real dataset that they choose. To ensure that students choose datasets that could lead to meaningful projects, we ask them to propose two possible datasets and explain how they fulfill criteria for the project (such as having enough variables to motivate at least 10 interesting graphs, a data structure complex enough to motivate at least one of the advanced topics in the second half of the semester, and an application nuanced enough to motivate at least three research questions). For their final project, teams must give a short oral presentation, submit a written report, and create an HTML document that is posted on our department website at stat.cmu.edu/capstoneresearch/. In this way, students not only demonstrate their ability to communicate statistical visualizations and analyses in several contexts, but also leave the class with a public-facing document that they can refer to. Indeed, we encourage students to link their projects on their resumes, such that they can showcase their data science portfolio to future employers and graduate programs. We also use these online documents to run a graphics contest, where faculty and PhD students vote on the best graphics of the semester, such that undergraduates can get external validation of their work.
@@
Our Master’s level course, titled Data Visualization, is an accelerated 7-week course that serves between 25-35 students in the first half of the fall semester. This course appears at the beginning of the Master of Science in Applied Data Science (MADS) curriculum, which spans two semesters and is targeted for students pursuing an industry position as a data scientist, data analyst, or data engineer. While there is similar content to the undergraduate course, there is a greater emphasis on computing and design, along with more advanced topics such as nonlinear dimension reduction techniques. An example of the course calendar and content is available at: https://ryurko.github.io/mads-36613-fall24/.

### v2

@@ 4 Conclusion and Discussion
First, we conducted a survey of data visualization courses taught at 135 top colleges and universities in the United States in the 2022-2023 academic year, in order to understand the landscape of data visualization courses. We identified 270 courses across many departments; the most common were related to computer science, data, information, and management, but departments in analytics, business, engineering, the humanities, journalism, public policy, statistics, the visual arts, and other fields were also represented. Thus, the instruction of data visualization is widespread and diverse. Most courses were not taught by statistics or data science departments, and thus understandably did not focus on statistical topics. Instead, most courses focused on visual storytelling, aesthetic design, dashboard design, specialized software, or other topics unique to their respective disciplines. It was especially uncommon for courses to focus on core aspects of statistical thinking: testing, uncertainty quantification, and modeling. Thus, statistics-related departments can provide a valuable perspective on data visualization that is unique to current course offerings.
@@
Second, we outlined three principles for teaching data visualization to encourage statistical thinking, using illustrative examples from our undergraduate data visualization course at Carnegie Mellon University. The first principle is that graphs should be complemented with inference. For example, the “statistical significance” of visual observations (e.g., differences in proportions or means) can be clarified with standard tests (e.g., chi-squared tests or ANOVA). This helps students understand how to combine graphs and analyses to make a cohesive argument, and encourages them to grapple with complex inferential issues (e.g., power, multiple testing, and visually overlapping confidence intervals). The second principle is that graphs are estimates and statistics that can be leveraged in an inferential framework. For example, graphs can have a null distribution that’s used in a testing procedure, and have uncertainty that can be quantified theoretically or via simulation with a bootstrap procedure. This helps students gain a deeper understanding of distributions and estimators by seeing how they operate with complex visuals beyond just univariate random variables. The third principle is that graphs can be used to motivate, teach, and interpret statistical analyses. Because visuals are a tried-and-true teaching tool, data visualization courses can be a natural conduit for students to engage with challenging concepts, like interactions in regression and principal component analysis. In this way, we’ve used data visualization as part of a spiral learning approach in our undergraduate curriculum, where students revisit topics several times (e.g., in a visualization context and in a theoretical context) as they master difficult material.
@@
That said, there are several limitations to our contributions, which motivate future work. Our survey is likely not a census of all data visualization courses, because (1) we used select keywords to identify potential courses, (2) we relied only on publicly available information to declare a course as a data visualization course, and (3) we only considered top colleges and universities according to the U.S. News and World Report in 2023. Furthermore, because our survey only considers courses taught in the 2022-2023 academic year, we have not considered how instruction on data visualization has possibly changed over time. Finally, we focused primarily on courses’ statistical content and did not form a full taxonomy of topics taught in data visualization courses. For example, we observed that many courses focused on visual storytelling and aesthetic design (particularly in business, journalism, the visual arts, and writing), but did not study which methods within storytelling and design these courses taught. Because different disciplines have their own specialized interest in data visualization, other fields may want to study which topics are covered in data visualization courses within their own discipline, perhaps using our publicly available dataset on these courses. Indeed, we hope our dataset allows others to explore the diversity of data visualization courses, including instructors and students seeking datasets for undergraduate projects.
@@ Acknowledgments
We would like to thank Peter Freeman, Alex Reinhart, and three anonymous reviewers for helpful comments that improved the presentation of this work.
@@
Here we discuss details about course design for our data visualization courses in the Department of Statistics and Data Science at Carnegie Mellon University. Specifically, we outline the structure and content for our undergraduate and master’s courses. Each course has its own intended audience and length. Thus, we hope these materials demonstrate how instructors in statistics-related fields can teach data visualization at the undergraduate and graduate level.
@@
Some readers may wonder how the set of students can be so diverse when the prerequisite is seemingly strict. One of our college’s general education requirements is that students must complete an introductory statistics course during their freshman year. Thus, students only have to take one additional statistics course in order to enroll in our course. Our reasoning for this prerequisite is to ensure that students have some familiarity with statistical inference, such that they can appreciate how inferential tools can complement data visualizations, which is one focus of the course.
Each week consists of two 50-minute lectures (one on Monday and one on Wednesday) and a 50-minute computer lab on Friday. Typically, 1-2 lectures are dedicated to ways to visualize particular data types (e.g., two-dimensional categorical data, one-dimensional quantitative data), and then 1-2 lectures are dedicated to inferential tools that can be used for that data type (e.g., chi-squared tests, Kolmogorov-Smirnov tests). Meanwhile, in the computer labs, students work through practice problems related to that week’s lecture content; along the way, they can ask each other and the teaching staff questions, such that the computer labs act as an active learning component. The typical week-by-week timeline and set of topics for the course is shown in Table 5. The first half of the semester focuses on visualizations and inferential tools involving one or two categorical and/or quantitative variables; this also corresponds to the taxonomy illustrated in Table 1. Meanwhile, the second half of the semester focuses on topics involving more complex data types: high-dimensional, spatial, longitudinal, and text.
@@
As for assessments, we have students submit weekly homework assignments, which involve creating visualizations and conducting analyses with real datasets based on the past week’s lecture and computer lab content. Furthermore, as a small participation grade, students are asked to critique a data visualization they found online or elsewhere once a month, in order to see how principles we discuss in class apply to real-world graphics. At midsemester, students complete a take-home midterm exam that assesses their ability to implement, navigate, and interpret the visualizations and inferential tools in Table 1. After the midterm exam, students are organized into teams of 3-4 people, and by the end of the semester complete a final project that consists of visualizations and analyses of a real dataset that they choose. To ensure that students choose datasets that could lead to meaningful projects, we ask them to propose two possible datasets and explain how they fulfill criteria for the project (such as having enough variables to motivate at least 10 interesting graphs, a data structure complex enough to motivate at least one of the advanced topics in the second half of the semester, and an application nuanced enough to motivate at least three research questions). For their final project, teams must give a short oral presentation, submit a written report, and create an HTML document that is posted on our department website at stat.cmu.edu/capstoneresearch/. In this way, students not only demonstrate their ability to communicate statistical visualizations and analyses in several contexts, but also leave the class with a public-facing document that they can reference. Indeed, we encourage students to link their projects on their resumes, such that they can showcase their data science portfolio to future employers and graduate programs. We also use these online documents to run a graphics contest, where faculty and PhD students vote on the best graphics of the semester, such that undergraduates can get external validation of their work.
@@
Our master’s level course, titled Data Visualization, is an accelerated 7-week course that serves approximately 30 students in the first half of the fall semester. This course appears at the beginning of our Master of Science in Applied Data Science (MADS) curriculum, which spans two semesters and is targeted for students pursuing an industry position as a data scientist, data analyst, or data engineer. While there is similar content to the undergraduate course, there is a greater emphasis on computing and design, along with more advanced topics such as nonlinear dimension reduction techniques. An example of the course calendar and content is available at: https://ryurko.github.io/mads-36613-fall24/. There, one can see how the first seven weeks of content in Table 5 is done in approximately two weeks, such that more time is spent on more advanced visualizations related to high-dimensional data, spatiotemporal data, text data, and interactive graphics with Shiny in R, as is appropriate for a graduate-level course. The website also includes the course’s lecture slides and coding demos for implementing visualizations and analyses in R, which may be helpful resources for instructors wanting to develop their own data visualization course.
Data Availability Statement
The data that support the findings of this study are openly available on the Open Science Framework (OSF) at https://osf.io/wdehn/?view_only=81a99ddcf51f492393ea567b3d5f37dc. They are also available in our replication materials at https://github.com/ryurko/teaching-data-viz.